# AMUN: Adversarial Machine UNlearning

### v1

@@ 1.1 Related Work
Early works in machine unlearning focused on exact solutions Cao and Yang (2015); Bourtoule et al. (2021); those ideas were adapted to unlearning in other domains such as graph neural networks Chen et al. (2022b) and recommendation systems Chen et al. (2022a). The extensive computational cost and utility loss resulted in the design of approximate methods. An example is the work of Ginart et al. (2019), who provide a definition of unlearning based on differential privacy. Works that followed sought solutions to satisfy those probabilistic guarantees (Ginart et al., 2019; Gupta et al., 2021; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021). However, the methods that satisfy these guarantees were only applied to simple models, such as $k$-means Ginart et al. (2019) , linear and logistic regression Guo et al. (2019); Izzo et al. (2021), convex-optimization problems Neel et al. (2021), or graph neural networks with no non-linearities Chien et al. (2022). Additional research was carried out to design more scalable approximate methods, those that can be applied to the models that are used in practice, including large neural networks Golatkar et al. (2020); Warnecke et al. (2021); Izzo et al. (2021); Thudi et al. (2022); Chen et al. (2023b); Liu et al. (2024); Fan et al. (2024). However, these approximate methods do not come with theoretical guarantees; their effectiveness are evaluated using membership inference attacks (MIAs). MIAs aim to determine whether a specific data sample was used in the training set of a trained model (Shokri et al., 2017; Yeom et al., 2018; Song et al., 2019; Hu et al., 2022; Carlini et al., 2022; Zarifzadeh et al., 2024), and is a common evaluation metric (Liu et al., 2024; Fan et al., 2024). s For further discussion on related works see Appendix A.
@@ 3.1 A Guiding Observation
Results: Figure 3 in Appendix D.1, shows the confidence scores (see § D.1 for details) for a ResNet-18 He et al. (2016) model that has been retrained on $\mathcal{D}-\mathcal{D_{\text{F}}}$, where $\mathcal{D}$ is the training set of CIFAR-10 Alex (2009) and the size of $\mathcal{D_{\text{F}}}$ (randomly chosen from $\mathcal{D}$) is $10\%$ and $50\%$ of the size of $\mathcal{D}$ (the first and second sub-figures, respectively).
@@ 3.2 Fine-tuning on Adversarial Examples
Results: As Figure 1 (and Figure 5 in Appendix D.3) shows, the test accuracy of the model does not deteriorate, even when it is being fine-tuned on only $\mathcal{D_{\text{A}}}$ (the dataset with wrong labels). See § 6.2.1 for further details.
@@ 4 Adversarial Machine UNlearning (AMUN)
The reason behind minimizing the distance of $\epsilon$ for each sample is to localize the changes to the decision boundary of the model as much as possible; this prevents changing the model’s behavior on other parts of the input space. For our experiments, we use PGD-50 (Madry, 2017) with $\ell_{2}$ norm bound as $\mathcal{A}_{\mathcal{F}}$. We set the step size of the gradient ascent in the attack to $0.1\times\epsilon$, which changes with the $\epsilon$ value. More details regarding the implementations of AMUN and prior unlearning methods and tuning their hyper-parameters can be found in Appendix B. Also, in Appendix D.5, we will show how using weaker attacks, such as Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014), might lead to lower performance of AMUN.
@@ 5 Evaluation Setup
In this section we elaborate on the details of evaluating different unlearning methods. More details (e.g., choosing the hyper-parameters) can be found in Appendix B.
@@ 5.1 Baseline Methods
We compare AMUN with FT (Warnecke et al., 2021), RL (Golatkar et al., 2020), GA (Thudi et al., 2022), BS (Chen et al., 2023b), $l_{1}$-Sparse (Liu et al., 2024), and SalUn (Fan et al., 2024). We also combine the weight saliency idea for masking the model parameters to limit the changes to the parameters during fine-tuning with AMUN and present its results as AMUN+SalUn (see Appendix C for more details). We use the same hyper-parameter tuning reported by prior works. For further details, see Appendix B.
@@ 6 Experiments
As a quick summary, our results show that: (1) AMUN effectively leads to unlearning the samples in $\mathcal{D_{\text{F}}}$: after unlearning $10\%$ of the samples of CIFAR-10 from a trained ResNet-18, RMIA cannot do better than random guessing (§ 6.1); (2) If we replace $\mathcal{D_{\text{A}}}$ with any of the aforementioned substitutes, the model’s accuracy significantly deteriorates, especially when there is no access to $\mathcal{D_{\text{R}}}$ (§ 6.2.1); (3) AMUN is as effective for unlearning on models that are adversarially robust (§ 6.2.2); (4) using weaker attack methods, such as FGSM, in AMUN hurts the effectiveness by not finding the adversarial examples that are very close to the samples in $\mathcal{D_{\text{F}}}$. However, they still outperform prior methods (Appendix D.5). The transferred adversarial examples are effective as well (Appendix D.4); and (5) AMUN outperforms other unlearning methods when handling multiple unlearning requests (§ 6.3).
@@ 6.1 Effectiveness of AMUN
We also present the results when MIS is used as the evaluation metric in Tables 7 and 8 in Appendix E, which similarly shows AMUN’s dominance in different unlearning settings. Moreover, we evaluate the combination of AMUN and SalUn (see Appendix C for details) and present its results as AMUNSalUn in these tables. AMUNSalUn slightly improves the results of AMUN in the setting where there is no access to $\mathcal{D_{\text{R}}}$, by filtering the parameters that are more relevant to $\mathcal{D_{\text{F}}}$ during fine-tuning.
@@ 6.2 Ablation Studies
In this subsection, we first elaborate on the effect of fine-tuning a model on its adversarial examples and compare it to the cases where either the samples or labels of this dataset change (§ 6.2.1). We then discuss AMUN’s efficacy on models that are already robust to adversarial examples (§ 6.2.2). We present other ablation studies on using weaker, but faster, adversarial attacks in Algorithm 1 (Appendix D.5). In Appendix D.4, we utilize transferred adversarial examples for unlearning, as this can expedite handling the unlearning from a newly trained model for which adversarial examples on similar architectures are available.
@@ 6.2.1 Fine-tuning on Adversarial Examples
Results: In Fig. 1, Adv, from the left sub-figure to the right sub-figure, shows the results when $\mathcal{D}\cup\mathcal{D_{\text{A}}}$, $\mathcal{D_{\text{F}}}\cup\mathcal{D_{\text{A}}}$, and $\mathcal{D_{\text{A}}}$ is used for fine-tuning the model, respectively. Orig, Adv-RS, Adv-RL, Orig-RL, and Orig-AdvL show the results when $\mathcal{D_{\text{A}}}$ for each of these sub-figures is replaced by $\mathcal{D_{\text{F}}}$, $\mathcal{D_{\text{A}}}_{RS}$, $\mathcal{D_{\text{A}}}_{RL}$, $\mathcal{D}_{RL}$, and $\mathcal{D}_{AdvL}$, respectively. As the figure shows, the specific use of adversarial examples with the mispredicted labels matters in keeping the model’s test accuracy, especially as we move from the leftmost sub-figure (having access to $\mathcal{D_{\text{R}}}$) to the right one (only using $\mathcal{D_{\text{A}}}$ or its substitutes). This is due to the fact that the samples in $\mathcal{D_{\text{A}}}$, in contrast to the other constructed datasets, belong to the natural distribution learned by the trained model. Therefore, even if we only fine-tune the ResNet-18 model on $\mathcal{D_{\text{A}}}$, we still do not lose much in terms of model’s accuracy on $\mathcal{D_{\text{T}}}$. This is a surprising observation, as $\mathcal{D_{\text{A}}}$ contains a set of samples with wrong predictions! Fig. 5 in Appendix D.3 shows similar results when size of $\mathcal{D_{\text{F}}}$ is $50\%$ of $|\mathcal{D}|$.
@@ 6.2.2 Adversarially Robust Models
We evaluate the effectiveness of AMUN when the trained model is adversarially robust. One of the most effective methods in designing robust models is adversarial training which targets smoothing the model’s prediction function around the training samples Salman et al. (2019). This has been shown to provably enhance the adversarial robustness of the model Cohen et al. (2019). However, this method is very costly and impractical for larger models. There is a separate line of work that try to achieve this smoothness by controlling the Lipschitz constant of the models Szegedy (2013). The method proposed by Boroojeny et al. (2024) is much faster than adversarial training and their results show a significant improvement in the robust accuracy. Since evaluations with RMIA require training 128 reference models, we use their clipping method to evaluate the effectiveness of AMUN for unlearning $10\%$ and $50\%$ of the samples from robust ResNet-18 models trained on CIFAR-10.
Results: As shown in Table 3, even when there is no access to $\mathcal{D_{\text{R}}}$, AMUN still results in effective unlearning for these models; RMIA does not do better than random guessing for predicting $\mathcal{D_{\text{F}}}$ from $\mathcal{D_{\text{T}}}$. Table 4 in Appendix D.2 shows similar results for the case with access to $\mathcal{D_{\text{R}}}$. As Fig. 3 (right) shows, in the robust models, more than $97\%$ of the adversarial examples are further away from their corresponding training samples, compared to this distance for the original models. However, this does not interfere with the performance of AMUN because these robust models are smoother and tend to be more regularized. This regularization, which prevents them from overfitting to the training samples is in fact a contributing factor to the improved generalization bounds for these models Bartlett et al. (2017). This in itself contributes to enhanced resilience against MIAs. As seen in Tables 3 and 4, even for the clipped models retrained on $\mathcal{D_{\text{R}}}$, the AUC score of RMIA for predicting $\mathcal{D_{\text{R}}}$ from $\mathcal{D_{\text{F}}}$ (FR AUC) is very low, which shows that these smoother models .
@@ 6.3 Continuous Unlearning
Results: Fig. 2 shows an overview of the results for both settings of unlearning (with or without access to $\mathcal{D_{\text{R}}}$). This figure shows the effectiveness of unlearning by depicting how the samples in $\mathcal{D_{\text{F}}}$ are more similar to the test samples ($\mathcal{D_{\text{T}}}$) rather than the remaining samples ($\mathcal{D_{\text{R}}}$). The value on the y-axis shows the difference of the area under the ROC curve (AUC) for predicting $\mathcal{D_{\text{R}}}$ from $\mathcal{D_{\text{F}}}$ and $\mathcal{D_{\text{F}}}$ from $\mathcal{D_{\text{T}}}$. For the plots of each of these values separately, see Appendix F. AMUN-A performs better than all the other unlearning methods for all the steps of unlearning. Although AMUN also outperforms all the prior unlearning methods, it slightly under-performs compared to AMUN-A. This is expected, as the model’s decision boundary slightly changes after each unlearning request and the adversarial examples generated for the original model might not be as effective as those ones generated for the new model. Note that for this experiment, we did not perform hyper-parameter tuning for any of the unlearning methods, and used the same ones derived for unlearning $10\%$ of the dataset presented in § 6.1. For further discussion of the results see Appendix F.

### v2

@@ 1.1 Related Work
Early works in machine unlearning focused on exact solutions Cao and Yang (2015); Bourtoule et al. (2021); those ideas were adapted to unlearning in other domains such as graph neural networks Chen et al. (2022b) and recommendation systems Chen et al. (2022a). The extensive computational cost and utility loss resulted in the design of approximate methods. An example is the work of Ginart et al. (2019), who provide a definition of unlearning based on differential privacy. Works that followed sought solutions to satisfy those probabilistic guarantees (Ginart et al., 2019; Gupta et al., 2021; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021). However, the methods that satisfy these guarantees were only applied to simple models, such as $k$-means Ginart et al. (2019) , linear and logistic regression Guo et al. (2019); Izzo et al. (2021), convex-optimization problems Neel et al. (2021), or graph neural networks with no non-linearities Chien et al. (2022). Additional research was carried out to design more scalable approximate methods, those that can be applied to the models that are used in practice, including large neural networks Golatkar et al. (2020); Warnecke et al. (2021); Izzo et al. (2021); Thudi et al. (2022); Chen et al. (2023b); Liu et al. (2024); Fan et al. (2024). However, these approximate methods do not come with theoretical guarantees; their effectiveness are evaluated using membership inference attacks (MIAs). MIAs aim to determine whether a specific data sample was used in the training set of a trained model (Shokri et al., 2017; Yeom et al., 2018; Song et al., 2019; Hu et al., 2022; Carlini et al., 2022; Zarifzadeh et al., 2024), and is a common evaluation metric (Liu et al., 2024; Fan et al., 2024). s For further discussion on related works see Appendix B.
@@ 3.1 A Guiding Observation
Results: Figure 3 in Appendix E.1, shows the confidence scores (see § E.1 for details) for a ResNet-18 He et al. (2016) model that has been retrained on $\mathcal{D}-\mathcal{D_{\text{F}}}$, where $\mathcal{D}$ is the training set of CIFAR-10 Alex (2009) and the size of $\mathcal{D_{\text{F}}}$ (randomly chosen from $\mathcal{D}$) is $10\%$ and $50\%$ of the size of $\mathcal{D}$ (the first and second sub-figures, respectively).
@@ 3.2 Fine-tuning on Adversarial Examples
Results: As Figure 1 (and Figure 5 in Appendix E.3) shows, the test accuracy of the model does not deteriorate, even when it is being fine-tuned on only $\mathcal{D_{\text{A}}}$ (the dataset with wrong labels). See § 6.2.1 for further details.
@@ 4 Adversarial Machine UNlearning (AMUN)
The reason behind minimizing the distance of $\epsilon$ for each sample is to localize the changes to the decision boundary of the model as much as possible; this prevents changing the model’s behavior on other parts of the input space. For our experiments, we use PGD-50 (Madry, 2017) with $\ell_{2}$ norm bound as $\mathcal{A}_{\mathcal{F}}$. We set the step size of the gradient ascent in the attack to $0.1\times\epsilon$, which changes with the $\epsilon$ value. More details regarding the implementations of AMUN and prior unlearning methods and tuning their hyper-parameters can be found in Appendix C. Also, in Appendix E.5, we will show how using weaker attacks, such as Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014), might lead to lower performance of AMUN.
@@
4.1 Influencing Factors
In this section, we derive an upper-bound on the $2$-norm of the difference of the parameters of the unlearned model and the retrained model (which are gold-standard for unlearning) that illuminated the influencing factors in the effectiveness of AMUN. To prove this theorem, we make assumptions that are common in the certified unlearning literature. The proof is given in Appendix A.
Theorem 4.1.
Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=\{1,\dots,N\}}$ be a dataset of $N$ samples and without loss of generality let $(x_{n},y_{n})$ (henceforth represented as $(x,y)$ for brevity) be the sample that needs to be forgotten and $(x^{\prime},y^{\prime})$ be its corresponding adversarial example used by AMUN such that $\|x-x^{\prime}\|_{2}=\delta$. Let ^$\hat{\mathcal{R}}(w)$ represent the (unnormalized) empirical loss on $\mathcal{D}^{\prime}=\mathcal{D}\cup\{(x^{\prime},y^{\prime})\}$ for a model $f$ that is parameterized with $w$. We assume that $f$ is $L$-Lipschitz with respect to the inputs and $\hat{\mathcal{R}}$ is $\beta$-smooth and convex with respect to the parameters. Let $\theta_{o}$ represent the parameters corresponding to the model originally trained on $\mathcal{D}$ and $\theta_{u}$ be the parameters derived when the model is trained on $\mathcal{D}-\{(x,y)\}$. We also assume that both the original and retrained models achieve near-00 loss on their training sets. After AMUN performs fine-tuning on $\mathcal{D}^{\prime}$ using one step of gradient descent with a learning rate of $\frac{1}{\beta}$ to derive parameters $\theta^{\prime}$, we get the following upper-bound for the distance of the unlearned model and the model retrained on $\mathcal{D}-\{(x,y)\}$ (gold standard of unlearning):
$\|\theta^{\prime}-\theta_{u}\|_{2}^{2}\leq\|\theta_{o}-\theta_{u}\|_{2}^{2}+\frac{2}{\beta}(L\delta-C),$
where $C=\ell(f_{\theta_{o}}(x^{\prime}),y)+\ell(f_{\theta^{\prime}}(x^{\prime}),y^{\prime})-\ell(f_{\theta_{u}}(x),y)-\ell(f_{\theta_{u}}(x^{\prime}),y^{\prime})$.
According to the bound in Theorem 4.1, a lower Lipschitz constant of the model ($L$) and adversarial examples that are closer to the original samples (lower value for $\delta$) lead to a smaller upper bound. A larger value of $C$ also leads to a improved upper-bound. In the following we investigate the factors that lead to a larger value for $C$, which further clarifies some of influencing factors in the effectiveness of AMUN:
  • Higher quality of adversarial example in increasing the loss for the correct label on the original model, which leads to larger value for $\ell(f_{\theta_{o}}(x^{\prime}),y)$.
  • Transferability of the adversarial example generated on the original model to the retrained model to decrease its loss for the wrong label, which leads to a lower value for $\ell(f_{\theta_{u}}(x^{\prime}),y^{\prime})$.
  • Early stopping and using appropriate learning rate during fine-tuning phase of unlearning to avoid overfitting to the adversarial example, which does not allow low values for $\ell(f_{\theta^{\prime}}(x^{\prime}),y^{\prime})$.
  • The generalization of the retrained model to the unseen samples, which leads to a lower value for $\ell(f_{\theta_{u}}(x),y)$.
Note that the first two implications rely on the strength of the adversarial example in addition to being close to the original sample. The second bullet, which relies on the transferability of adversarial examples, has been shown to improve as the Lipschitz constant decreases (Ebrahimpour-Boroojeny et al., 2024). The third bullet point is a natural implication which also holds for other unlearning methods that rely on the fine-tuning of the model. The fourth bullet point is not relevant to the unlearning method and instead relies on the fact that the retrained model should have good generalizability to unseen samples; it implies that as the size of $\mathcal{D_{\text{F}}}$ increases (i.e., $|\mathcal{D_{\text{R}}}|$ decreases) and the performance of the retrained model decreases, the effectiveness of the unlearning model also decreases. This is also intuitively expected in the unlearning process. Hence, the proved theorem also justifies our earlier intuitions about the need for good quality adversarial examples that are as close as possible to the original samples (which is the goal of Algorithm 1).
@@ 5 Evaluation Setup
In this section we elaborate on the details of evaluating different unlearning methods. More details (e.g., choosing the hyper-parameters) can be found in Appendix C.
@@ 5.1 Baseline Methods
We compare AMUN with FT (Warnecke et al., 2021), RL (Golatkar et al., 2020), GA (Thudi et al., 2022), BS (Chen et al., 2023b), $l_{1}$-Sparse (Liu et al., 2024), and SalUn (Fan et al., 2024). We also combine the weight saliency idea for masking the model parameters to limit the changes to the parameters during fine-tuning with AMUN and present its results as AMUN+SalUn (see Appendix D for more details). We use the same hyper-parameter tuning reported by prior works. For further details, see Appendix C.
@@ 6 Experiments
As a quick summary, our results show that: (1) AMUN effectively leads to unlearning the samples in $\mathcal{D_{\text{F}}}$: after unlearning $10\%$ of the samples of CIFAR-10 from a trained ResNet-18, RMIA cannot do better than random guessing (§ 6.1); (2) If we replace $\mathcal{D_{\text{A}}}$ with any of the aforementioned substitutes, the model’s accuracy significantly deteriorates, especially when there is no access to $\mathcal{D_{\text{R}}}$ (§ 6.2.1); (3) AMUN is as effective for unlearning on models that are adversarially robust (§ 6.2.2); (4) using weaker attack methods, such as FGSM, in AMUN hurts the effectiveness by not finding the adversarial examples that are very close to the samples in $\mathcal{D_{\text{F}}}$. However, they still outperform prior methods (Appendix E.5). The transferred adversarial examples are effective as well (Appendix E.4); and (5) AMUN outperforms other unlearning methods when handling multiple unlearning requests (§ 6.3).
@@ 6.1 Effectiveness of AMUN
We also present the results when MIS is used as the evaluation metric in Tables 8 and 9 in Appendix F, which similarly shows AMUN’s dominance in different unlearning settings. Moreover, we evaluate the combination of AMUN and SalUn (see Appendix D for details) and present its results as AMUNSalUn in these tables. AMUNSalUn slightly improves the results of AMUN in the setting where there is no access to $\mathcal{D_{\text{R}}}$, by filtering the parameters that are more relevant to $\mathcal{D_{\text{F}}}$ during fine-tuning.
@@ 6.2 Ablation Studies
In this subsection, we first elaborate on the effect of fine-tuning a model on its adversarial examples and compare it to the cases where either the samples or labels of this dataset change (§ 6.2.1). We then discuss AMUN’s efficacy on models that are already robust to adversarial examples (§ 6.2.2). We present other ablation studies on using weaker, but faster, adversarial attacks in Algorithm 1 (Appendix E.5). In Appendix E.4, we utilize transferred adversarial examples for unlearning, as this can expedite handling the unlearning from a newly trained model for which adversarial examples on similar architectures are available.
@@ 6.2.1 Fine-tuning on Adversarial Examples
Results: In Fig. 1, Adv, from the left sub-figure to the right sub-figure, shows the results when $\mathcal{D}\cup\mathcal{D_{\text{A}}}$, $\mathcal{D_{\text{F}}}\cup\mathcal{D_{\text{A}}}$, and $\mathcal{D_{\text{A}}}$ is used for fine-tuning the model, respectively. Orig, Adv-RS, Adv-RL, Orig-RL, and Orig-AdvL show the results when $\mathcal{D_{\text{A}}}$ for each of these sub-figures is replaced by $\mathcal{D_{\text{F}}}$, $\mathcal{D_{\text{A}}}_{RS}$, $\mathcal{D_{\text{A}}}_{RL}$, $\mathcal{D}_{RL}$, and $\mathcal{D}_{AdvL}$, respectively. As the figure shows, the specific use of adversarial examples with the mispredicted labels matters in keeping the model’s test accuracy, especially as we move from the leftmost sub-figure (having access to $\mathcal{D_{\text{R}}}$) to the right one (only using $\mathcal{D_{\text{A}}}$ or its substitutes). This is due to the fact that the samples in $\mathcal{D_{\text{A}}}$, in contrast to the other constructed datasets, belong to the natural distribution learned by the trained model. Therefore, even if we only fine-tune the ResNet-18 model on $\mathcal{D_{\text{A}}}$, we still do not lose much in terms of model’s accuracy on $\mathcal{D_{\text{T}}}$. This is a surprising observation, as $\mathcal{D_{\text{A}}}$ contains a set of samples with wrong predictions! Fig. 5 in Appendix E.3 shows similar results when size of $\mathcal{D_{\text{F}}}$ is $50\%$ of $|\mathcal{D}|$.
@@ 6.2.2 Adversarially Robust Models
We evaluate the effectiveness of AMUN when the trained model is adversarially robust. One of the most effective methods in designing robust models is adversarial training which targets smoothing the model’s prediction function around the training samples Salman et al. (2019). This has been shown to provably enhance the adversarial robustness of the model Cohen et al. (2019). One of the effective adversarial training methods is by using TRADES loss introduced by (Zhang et al., 2019). We will use adversarially trained ResNet-18 models for unlearning $10\%$ of the samples in CIFAR-10. In addition, we will use another defense mechanism that is less costly and more practical for larger models. There is a separate line of work that try to achieve the same smoothness in model’s prediction boundary by controlling the Lipschitz constant of the models Szegedy (2013). The method proposed by Boroojeny et al. (2024) is much faster than adversarial training and their results show a significant improvement in the robust accuracy. We use their clipping method to evaluate the effectiveness of AMUN for unlearning $10\%$ and $50\%$ of the samples from robust ResNet-18 models trained on CIFAR-10.
Results: Table 5 in Appendix E.2 shows the results for the adversarially trained models. For the models with controlled Lipschitz continuity, the results are shows in Table 3 (no access to $\mathcal{D_{\text{R}}}$) and Table 4 in Appendix E.2 (with access to $\mathcal{D_{\text{R}}}$). As the results show, even when there is no access to $\mathcal{D_{\text{R}}}$, AMUN still results in effective unlearning for adversarially robust models; RMIA does not do better than random guessing for predicting $\mathcal{D_{\text{F}}}$ from $\mathcal{D_{\text{T}}}$. As Fig. 3 (right) shows, in the robust models, more than $97\%$ of the adversarial examples are further away from their corresponding training samples, compared to this distance for the original models. However, this does not interfere with the performance of AMUN because these robust models are smoother and tend to be more regularized. This regularization, which prevents them from overfitting to the training samples is in fact a contributing factor to the improved generalization bounds for these models Bartlett et al. (2017). This in itself contributes to enhanced resilience against MIAs. As seen in Tables 3 and 4, even for the clipped models retrained on $\mathcal{D_{\text{R}}}$, the AUC score of RMIA for predicting $\mathcal{D_{\text{R}}}$ from $\mathcal{D_{\text{F}}}$ (FR AUC) is very low, which shows that these smoother models .
@@ 6.3 Continuous Unlearning
Results: Fig. 2 shows an overview of the results for both settings of unlearning (with or without access to $\mathcal{D_{\text{R}}}$). This figure shows the effectiveness of unlearning by depicting how the samples in $\mathcal{D_{\text{F}}}$ are more similar to the test samples ($\mathcal{D_{\text{T}}}$) rather than the remaining samples ($\mathcal{D_{\text{R}}}$). The value on the y-axis shows the difference of the area under the ROC curve (AUC) for predicting $\mathcal{D_{\text{R}}}$ from $\mathcal{D_{\text{F}}}$ and $\mathcal{D_{\text{F}}}$ from $\mathcal{D_{\text{T}}}$. For the plots of each of these values separately, see Appendix G. AMUN-A performs better than all the other unlearning methods for all the steps of unlearning. Although AMUN also outperforms all the prior unlearning methods, it slightly under-performs compared to AMUN-A. This is expected, as the model’s decision boundary slightly changes after each unlearning request and the adversarial examples generated for the original model might not be as effective as those ones generated for the new model. Note that for this experiment, we did not perform hyper-parameter tuning for any of the unlearning methods, and used the same ones derived for unlearning $10\%$ of the dataset presented in § 6.1. For further discussion of the results see Appendix G.